OpenStudy (anonymous):

This year, Ms. Makratzaki's small hardware store had sales of $536,000. Five years ago, the store had $441,000 in sales, and last year, $517,000 in sales. Assuming a constant rate of increase, predict the amount of sales the store will have after 8 years.

OpenStudy (anonymous):

hey are u doin home work

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

see this year sales is $ 536 last year it was $ 517 that is difference of $ 19 and five years ago it was $ 441 Difference = $ 536 - $ 441 = $ 95 = $ 19 * 5 it means that every year sales increases by $ 19 so after 8 years it would be ???????

OpenStudy (anonymous):

u got the answer

OpenStudy (anonymous):

yah thanks

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

well ill probably feel stupid saying the wrong awnser but i got..........764?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

seriously its wrong how did i do it wrong?

OpenStudy (anonymous):

it is 536 + 19*8 =

OpenStudy (anonymous):

4440.....?

OpenStudy (anonymous):

oh no 536000 + 19000*8 = 536000 + 152000 = 688000

OpenStudy (anonymous):

thats the awnser?